Weston Corporation just paid a dividend of $3 a share (i.e., D0 = $3). The dividend is expected to grow 10% a year for the next 3 years and then at 3% a year thereafter. What is the expected dividend per share for each of the next 5 years? Do not round intermediate calculations. Round your answers to the nearest cent.

D1 = $

D2 = $

D3 = $

D4 = $

D5 = $
Answer #1
D1 (3*1.1)=$3.3
D2 (3.3*1.1)=$3.63
D3 (3.63*1.1)=$3.99(Approx).
D4 (3.99*1.03)=$4.11(Approx).
D5 (4.11*1.03)=$4.24(Approx).